OpenStudy (anonymous):

How can i do like 16!/4!12! ??? and can teach the rule ?

Parth (parthkohli):

This is a combination. 16c12 to be exact.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 16c12 = {16! \over (16 - 12)! \times 12!} }\)

OpenStudy (anonymous):

then it will 1/12!x12! ????

Parth (parthkohli):

Umm no

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {16 \times 15 \times 14 \times 13 \over4!} }\)

Parth (parthkohli):

I have cancelled 12! out of the scenario

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {16 \times 15 \times 14 \times 13 \over 4 \times 3 \times 2} }\)

OpenStudy (anonymous):

question if 4! then 4x3x2????

OpenStudy (anonymous):

and if 16! 16x15x14x13x12x11x10....x2 like this???

OpenStudy (phi):

yes. n! = n*(n-1)*(n-2)*...*1 so 4!= 4*3*2*1 to evaluate your 16!/(4! 12!) expand and then divide out